OpenStudy (anonymous):

find the exact radian value of arccsc(-2)

OpenStudy (kc_kennylau):

\[\theta=\csc^{-1}(-2)\]\[\csc\theta=-2\]\[\sin\theta=-0.5\]

OpenStudy (mathmale):

Great. Now, Amazinglysuperb, in which quadrant or quadrants could the angle (theta) lie if the sine of this (or these) angle(s) is negative?